154 mL of 3.87 mol/L barium hydroxide is needed to fully react with 146 mL of lithium sulfate solution what is the molar concentration of lithium sulfate solution


1
Expert's answer
2021-04-08T12:09:00-0400

Solution:

Balanced chemical equation:

Ba(OH)2 + Li2SO4 → BaSO4 + 2LiOH

According to the equation above: n(Ba(OH)2) = n(Li2SO4)

C(Ba(OH)2) × V(Ba(OH)2) = C(Li2SO4) × V(Li2SO4)

Divide to isolate C(Li2SO4):

C(Li2SO4) = C(Ba(OH)2) × V(Ba(OH)2) / V(Li2SO4)

Plug in the numbers and solve for C(Li2SO4):

C(Li2SO4) = (3.87 mol/L × 154 mL) / (146 mL) = 4.08 mol/L

C(Li2SO4) = 4.08 mol/L


Answer: The molar concentration of lithium sulfate (Li2SO4) solution is 4.08 mol/L.